80ml of medicine was made from a mix of three parts of water to one part solution. how much solution was used?
Answer by lwsshak3(11628)

80ml of medicine was made from a mix of three parts of water to one part solution. how much solution was used?
***
let x=amt of solution used
80-x=amt of water used
(80-x)/x=3
3x=80-x
4x=80
x=20
amt of solution used=20 ml